Two point charges lie on the x axis. A charge of + 2.60 pC is at the origin, and a charge of − 4.60 pC is at x=−10.0cm. What third charge should be placed at x=+24cm so that the total electric field at x=+12.0cm is zero?
Answer #1

Net field at P,

E = E1 + E2 - E3

Kq1/r1^2 + kq2/r2^2 - kq3/r3^2 = 0

2.6/10^2 - 4.6/24^2 - q3/12^2 = 0

Charge, q3 = 2.59 pC

Comment in case any doubt please rate my answer....